# Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior

### v4

@@ II Model and Preliminaries
In this section, we formulate the CP shaving game model and introduce its definitions. We begin with a two-agent (x and y), two-period (1 and 2) model, where the two agents can represent clusters of customers with different baseline demand conditions, and the two periods correspond to a peak period and an off-peak period, respectively. We denote the agent x’s baseline demand in periods 1 and 2 as $X_{1}$ and $X_{2}$, respectively, with $x$ representing its demand shifting. Similarly, we denote the agent y’s baseline demand as $Y_{1}$ and $Y_{2}$ with $y$ as its demand shifting. The system’s baseline demand, given by summing both agents’ baseline demands, is denoted as $S_{\mathrm{b},1},S_{\mathrm{b},2}$, while the system’s total demand, which includes both agents’ demand shifting, is denoted as $S_{1},S_{2}$, For simplicity, we often omit the $x,y$ when there is no ambiguity. Under this two-period setting, we define the direction of demand shifting as follows: shifting demand to period 1 (away from period 2) is considered positive, while shifting demand to period 2 (away from period 1) is considered negative, e.g., for agent x is $X_{1}+x$ and $X_{2}-x$. Depending on the agents’ demand-shifting behaviors, the roles of the peak and off-peak periods can interchange, i.e., either period 1 or period 2 can become the peak or off-peak period. We use $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}\in\mathbb{R}^{+}$ to denote the penalty parameter for agent x and y, representing the comfort loss or perceived cost of shifting demand. These parameters can be identified from shifting behavior suggested by historical data. The fixed CP charge price is denoted by $\pi\in\mathbb{R}^{+}$ and $I(x,y)$ is the step function (indicator function). Both agent gaming to reduce CP demand while maximizing their payoff, and the game model is defined as $G=(N,\mathcal{X},U)$, where
@@
  • $\mathcal{X}=\mathcal{X}_{\mathrm{x}}\times\mathcal{X}_{\mathrm{y}}$ is the strategy set formed by the product topology of each agent’s strategy set, where $\times$ is the topology product;
@@
During the game, agent x chooses strategy $x\in\mathcal{X}_{\mathrm{x}}$ to maximize its payoff $f_{\mathrm{x}}(x,y)$ with $y\in\mathcal{X}_{\mathrm{y}}$, and the payoff is to minimize its two periods’ costs,
$\max_{x}$ $f_{\mathrm{x}}(x,y)=-\pi(X_{1}+x)I(S_{1}(x,y)-S_{2}(x,y))-\pi(X_{2}-x)I(S_{2}(x,y)-S_{1}(x,y))-\alpha_{\mathrm{x}}x^{2},$ (1a)
$I(x)=\begin{cases}1&x\geq 0\\ 0&x<0\end{cases},$ (1b)
@@
Agent y’s payoff function $f_{\mathrm{y}}(x,y)$ follows the same structure by changing baseline demand to $Y_{1},Y_{2}$, demand shifting to $y$, and penalty parameter to $\alpha_{\mathrm{y}}$. Note that there is always a CP period even if two periods of demand are equal; here, we treat period 1 as the CP period, reflected as the equal sign of the step function (1b). Also, we focus on demand shifting and assume a constant energy rate besides the peak demand charge, so we don’t consider the energy cost.
It is obvious that the game is compact, convex, and bounded, where compact means each $\mathcal{X}_{\mathrm{x}},\mathcal{X}_{\mathrm{y}}$ is compact, convex means each $\mathcal{X}_{\mathrm{x}},\mathcal{X}_{\mathrm{y}}$ is convex, and bounded means each $f_{\mathrm{x}},f_{\mathrm{y}}$ is bounded. Note that we use a quadratic penalty to express a soft constraint to make the strategy set compact, and we have the linear correlation of the number of variables and agents. We then introduce the definitions of equilibrium and continuity.
@@
We define the following: (1) Pure-strategy Nash equilibrium (Nash [Nash Jr, 1950]). $(x^{*},y^{*})\in\mathcal{X}$ is a NE in pure strategies of the game $G$ if and only if (iff) $f_{\mathrm{x}}(x^{*},y^{*})\geq f_{\mathrm{x}}(x,y^{*})$ and $f_{\mathrm{y}}(x^{*},y^{*})\geq f_{\mathrm{y}}(x^{*},y)$ for every $x\in\mathcal{X}_{\mathrm{x}},y\in\mathcal{X}_{\mathrm{y}}$.
(2) Upper semi-continuity (u.s.c.) and lower semi-continuity (l.s.c.). The function $f:\mathcal{X}\rightarrow\mathbb{R}$ is called upper semi-continuous (u.s.c.) or lower semi-continuous (l.s.c.) if for every $x_{0}$ such that
@@
The step function makes the game analysis non-trivial. We thus separate the overall system into two periods corresponding to CP period in 1 and 2, i.e., $\mathcal{X}_{\mathrm{cp}1}=\{x,y|S_{1}(x,y)\geq S_{2}(x,y)\}$ and $\mathcal{X}_{\mathrm{cp}2}=\{x,y|S_{1}(x,y)<S_{2}(x,y)\}$, with the payoff functions, take agent x for example,
@@
It is obvious that each payoff function is concave, and by applying the first-order optimality condition [Boyd and Vandenberghe, 2004], we have
@@
    Critical point. $r_{\mathrm{x}}$ is the critical point of the payoff function $f_{\mathrm{x},2}$ described in (3), and the critical point of the payoff function $f_{\mathrm{x},1}$ is $-r_{\mathrm{x}}$. $r_{\mathrm{y}}$ follows the same idea.
    $r_{\mathrm{x}}=\frac{\pi}{2\alpha_{\mathrm{x}}},r_{\mathrm{y}}=\frac{\pi}{2\alpha_{\mathrm{y}}},$ (4a)
@@
    Agent balance point. $b_{\mathrm{x}},b_{\mathrm{y}}$ are the demand difference of agent x and y in two period periods
    $b_{\mathrm{x}}=\frac{X_{2}-X_{1}}{2},b_{\mathrm{y}}=\frac{Y_{2}-Y_{1}}{2},$ (4b)
@@
    System balance point. $b$ is the demand difference of the overall system in two period periods
    $b=b_{\mathrm{x}}+b_{\mathrm{y}}=\frac{S_{\mathrm{b},2}-S_{\mathrm{b},1}}{2}.$ (4c)
@@
    System average demand. $S$ is the average baseline demand of the overall system in two period periods is
@@
The critical point indicates the maximum demand each agent can shift to avoid the CP charge within one period. The agent balance allows each agent to shave its demand flat in two periods, which is the economic maximum demand shifting in the entire system. The reason is that once the agent balance point is reached, the CP charge is constant (half of the demand) regardless of the opponent’s strategy. Thus, $\min\{r_{\mathrm{x}},b_{\mathrm{x}}\}$ or $\max\{-r_{\mathrm{x}},b_{\mathrm{x}}\}$ define the maximum shifting capability of agent x. The system average demand is useful to show the CP charge when the system demand is balanced in the two period periods. We then define the capable and non-capable agents.
@@
Capable and non-capable agents. According to Definition 2, given agent x with baseline demand $X_{1},X_{2}$, CP charge $\pi$, and penalty parameter $\alpha_{\mathrm{x}}$, the agent x is capable if it satisfy
$-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},$ (5)
Define the agent that didn’t satisfy (5) as non-capable agent, including upper non-capable agent with $b_{\mathrm{x}}>r_{\mathrm{x}}$ and lower non-capable agent with $b_{\mathrm{x}}<-r_{\mathrm{x}}$.
According to Definition 2, (5) means agent x is economically capable of balancing its demand in the two periods during the game $G$ to reduce the CP charge. The reason is that the agent’s balance point $b_{\mathrm{x}}$ is within their critical point $[-r_{\mathrm{x}},r_{\mathrm{x}}]$ of both periods, indicating that they will reach the balance point before the critical point when shifting demand.
III Equilibrium analysis of the coincident peak game
In this section, we first show the properties of the two-agent two-period CP game $G$ with the specific agent type, and prove the pure-strategy NE of the CP game $G$.
@@ III-A Game properties
We first show the property of CP game $G$. Determined by the parameters of all agents, the game performs differently.
@@
Concave, quasiconcave/discontinuous, and non-concave/discontinuous CP game. Given critical point $r_{\mathrm{x}},r_{\mathrm{y}}$ and balance point $b_{\mathrm{x}},b_{\mathrm{y}},b$ as define in Definition 2, the two-agent two-period CP game $G$ satisfies one of the following:
@@
    Concave CP game. $G$ is concave that all agents’ payoff function $f_{\mathrm{x}}(x,y)$ ($f_{\mathrm{y}}(x,y)$) is concave in $x\in\mathcal{X}_{\mathrm{x}}$ ($y\in\mathcal{X}_{\mathrm{y}}$) for each $y\in\mathcal{X}_{\mathrm{y}}$ ($x\in\mathcal{X}_{\mathrm{x}}$) under the conditions of
    $b$ $<-r_{\mathrm{x}}-r_{\mathrm{y}},\text{if }S_{\mathrm{b},1}\geq S_{\mathrm{b},2},$ (6a)
    $b$ $>r_{\mathrm{x}}+r_{\mathrm{y}},\text{if }S_{\mathrm{b},1}<S_{\mathrm{b},2},$ (6b)
@@
    Quasiconcave CP game. $G$ is quasiconcave/discontinuous that all agents’ payoff function $f_{\mathrm{x}}(x,y)$ ($f_{\mathrm{y}}(x,y)$) is quasiconcave in $x\in\mathcal{X}_{\mathrm{x}}$ ($y\in\mathcal{X}_{\mathrm{y}}$) for each $y\in\mathcal{X}_{\mathrm{y}}$ ($x\in\mathcal{X}_{\mathrm{x}}$[Reny, 1999] under the conditions of
    $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},\ -r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}};$ (6c)
@@
    Non-concave CP game. $G$ is non-concave/discontinuous if it is not concave or quasiconcave. The condition is
    $\{-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}\}\cap$ [$[\{b_{\mathrm{x}}<-r_{\mathrm{x}}\}\cup\{b_{\mathrm{x}}>r_{\mathrm{x}}\}\cup\{b_{\mathrm{y}}<-r_{\mathrm{y}}\}\cup\{b_{\mathrm{y}}>r_{\mathrm{y}}\}].$ (6d)
@@
The payoff function is quadratic if only one possible CP period during the gaming process, indicating a concave and continuous game. The indicator function discontinues the game when the CP period changes. According to the definition of the quasiconcave function, we separate three cases: 1) Agent x can switch the CP period before reaching their critical points in both periods 1 and 2; 2) Agent x can reach the critical point in period 2 before switching, and the function is l.s.c. as defined in Definition 1; 3) Agent x can reach the critical point in period 1 before switching, and the function is u.s.c. as defined in Definition 1. We show only the first case is able for both agents’ functions to be quasiconcave. Then the non-concave condition is derived from the complementary set of concave and quasiconcave conditions. Also, note that if one agent’s payoff function is quasiconcave or non-concave, the payoff functions of the other agents must also be quasiconcave or non-concave. The detailed proof is provided in the appendix. ∎
This Proposition shows the game’s properties depend on the relationship between the agent’s critical point and balance point, affected by baseline demand $X_{1},X_{2},Y_{1},Y_{2}$, CP charge parameters $\pi$, and shifting penalty parameters $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$. In the concave game conditions, all agents can’t change the CP period together, and only one period is active, determined by system baseline conditions $S_{\mathrm{b},1}$ and $S_{\mathrm{b},2}$. If all agents are capable according to Definition 3, the CP game is quasiconcave. Otherwise, when both agents can change the CP period together, but one is non-capable, the CP game is non-concave. In practical systems, due to the variant agents’ shifting capability, all three types of games are likely to appear. Empirical studies on demand response indicate that some industrial consumers with automation can shift demand efficiently, while some are inefficient in shifting [Siano, 2014]. With these properties of CP game $G$, we then analyze its NE.

### v5

@@ II Model and Preliminaries
In this section, we formulate the CP shaving game model and introduce some definitions. We begin with a two-agent (x and y), two-period (1 and 2) model, where the two agents can represent clusters of customers with different baseline demand conditions, and the two periods correspond to a CP period and an off-CP period, respectively. We let agent x’s baseline demand in periods 1 and 2 be $X_{1}$ and $X_{2}$, respectively. Similarly, agent y’s baseline demand in periods 1 and 2 are, respectively, $Y_{1}$ and $Y_{2}$. The system’s baseline demand in periods 1 and 2, obtained by summing both agents’ baseline demands in that period, are denoted $S_{\mathrm{b},1}$ and $S_{\mathrm{b},2}$ respectively. Without loss of generality, we assume the baseline demand in period 2 is greater than in period 1, i.e., $S_{\mathrm{b},2}>S_{\mathrm{b},1}$. In the event that is not true for the given instance, we can interchange the roles of the two periods, and all the analysis will follow. Agents have the flexibility to shift some of their consumption from one period to the other. We let $x$ and $y$ denote, respectively, the amount of demand shifted by agents x and y, respectively, from period 2 to period 1. In other words, for any given $x$, the demand of agent x in the two periods are, respectively, $X_{1}+x$ and $X_{2}-x$; similarly for $y$ and agent y. Note that, without loss of generality, we further assume that agent x’s baseline demand in period 2 is greater than in period 1. i.e., $X_{2}>X_{1}$. This ensures that agent x’s demand shifting $x$ is non-negative at the equilibrium, while agent y’s shifting $y$ remains unrestricted. Still, the roles of agents x and y are interchangeable, so all results remain valid. The system’s total demand, which includes both agents’ demand shifting, is denoted as $S_{1}$ and $S_{2}$ in periods 1 and 2, respectively. Depending on the agents’ demand-shifting behaviors, the roles of the CP and off-CP periods can change, i.e., either period 1 or period 2 can become the CP or off-CP period. We use $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}\in\mathbb{R}^{+}$ to denote the penalty parameter for agent x and y, representing the comfort loss or perceived cost of shifting demand. These parameters can be identified from shifting behavior suggested by historical data. The fixed CP price is denoted by $\pi\in\mathbb{R}^{+}$ and $I(x,y)$ is the step function (indicator function). The agents choose $x$ and $y$ to maximize their payoffs. Formally, the game model is defined as $G=(N,\mathcal{X},U)$, where
@@
  • $\mathcal{X}=\mathcal{X}_{\mathrm{x}}\times\mathcal{X}_{\mathrm{y}}$ is the strategy set formed by the product topology of each agent’s strategy set;
@@
During the game, agent x chooses strategy $x\in\mathcal{X}_{\mathrm{x}}$ to maximize its payoff $f_{\mathrm{x}}(x,y)$ with $y\in\mathcal{X}_{\mathrm{y}}$ chosen by agent y. The payoff to agent x is simply the negative of its costs over the two time periods, and is given by:
$\max_{x}$ $f_{\mathrm{x}}(x,y)=-\pi(X_{1}+x)I(x,y)-\pi(X_{2}-x)(1-I(x,y))-\alpha_{\mathrm{x}}x^{2},$ (1a)
$I(x,y)=\begin{cases}1&S_{1}(x,y)-S_{2}(x,y)\geq 0\\ 0&S_{1}(x,y)-S_{2}(x,y)<0\end{cases},$ (1b)
@@
Agent y’s payoff function $f_{\mathrm{y}}(x,y)$ follows the same structure by changing baseline demand to $Y_{1},Y_{2}$, demand shifting to $y$, and penalty parameter to $\alpha_{\mathrm{y}}$. Note that there is always a CP period (and a corresponding charge) even if demands in the two periods are equal; in that case, we treat period 1 as the CP period, as reflected in our definition of the step function $I(\cdot)$ (27b). Also, we focus on demand shifting and assume a constant energy rate besides the peak demand charge, so we do not consider the energy cost.
We next introduce the definitions of equilibrium and continuity.
@@
We define the following: (1) Pure-strategy Nash equilibrium ([Nash Jr,, 1950]). $(x^{*},y^{*})\in\mathcal{X}$ is a NE in pure strategies of the game $G$ if and only if (iff) $f_{\mathrm{x}}(x^{*},y^{*})\geq f_{\mathrm{x}}(x,y^{*})$ and $f_{\mathrm{y}}(x^{*},y^{*})\geq f_{\mathrm{y}}(x^{*},y)$ for every $x\in\mathcal{X}_{\mathrm{x}},y\in\mathcal{X}_{\mathrm{y}}$.
(2) Upper semi-continuity (u.s.c.) and lower semi-continuity (l.s.c.). The function $f:\mathcal{X}\rightarrow\mathbb{R}$ is called u.s.c. or l.s.c. if for every $x_{0}$ such that
@@
The step function makes the game analysis non-trivial. We separate the overall system into two periods corresponding to CP period in 1 and 2, i.e., $\mathcal{X}_{\mathrm{cp}1}=\{x,y|S_{1}(x,y)\geq S_{2}(x,y)\}$ and $\mathcal{X}_{\mathrm{cp}2}=\{x,y|S_{1}(x,y)<S_{2}(x,y)\}$, with the payoff functions, take agent x for example,
@@
Note that each payoff function is concave, and by applying the first-order optimality condition [Boyd and Vandenberghe,, 2004], we have
@@
    Critical points for agents x and y:
    $r_{\mathrm{x}}=\frac{\pi}{2\alpha_{\mathrm{x}}},\quad r_{\mathrm{y}}=\frac{\pi}{2\alpha_{\mathrm{y}}}.$ (4a)
@@
    Agent balance points for agents x and y:
    $b_{\mathrm{x}}=\frac{X_{2}-X_{1}}{2},\quad b_{\mathrm{y}}=\frac{Y_{2}-Y_{1}}{2}.$ (4b)
@@
    System balance point:
    $b=\frac{S_{\mathrm{b},2}-S_{\mathrm{b},1}}{2}.$ (4c)
@@
    System average demand:
@@
The critical points $r_{\mathrm{x}},r_{\mathrm{y}}$ are determined by the payoff function in period 2, as exemplified in $f_{\mathrm{x},2}$ (3). These points represent the optimal demand shifting for each agent within period 2 to maximize its payoff. Naturally, the corresponding optimal shifting amounts for period 1 are $-r_{\mathrm{x}},-r_{\mathrm{y}}$. The agent balance points $b_{\mathrm{x}},b_{\mathrm{y}}$ represent inter-period demand shifting. For instance, if agent x chooses to shift $b_{x}$ from period 2 to period 1, its demand in both periods will be identical, and the same applies to agent y. Furthermore, if the total shifted demand across both agents equals $b$, then the system demand in periods 1 and 2 will be balanced and equal to $S$, which is the system average demand. Given the assumption that $S_{\mathrm{b},1}<S_{\mathrm{b},2}$ and $X_{1}<X_{2}$, it follows that $b_{\mathrm{x}}>0$ and $b>0$.
We then define capable and non-capable agents using agent x as an example, with agent y following the same definition.
@@
Capable and non-capable agents. According to Definition 2, given agent x with baseline demand $X_{1},X_{2}$, CP price $\pi$, and penalty parameter $\alpha_{\mathrm{x}}$, we define the agent x is capable if it satisfies
$-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}.$ (5)
Define an agent who does not satisfy (5) as a non-capable agent, including upper non-capable agent with $b_{\mathrm{x}}>r_{\mathrm{x}}$ and lower non-capable agent with $b_{\mathrm{x}}<-r_{\mathrm{x}}$.
According to Definition 2, (5) means agent x is economically capable of balancing its demand in the two periods to reduce the CP charge. This is because demand shifting will first reach the balance point before reaching the critical point. If agent x cannot balance its demand across two periods, its maximum shifting capability is constrained by the optimal shifting amount within a single period. Thus, the maximum shifting capability of agent x is given by $\min\{r_{\mathrm{x}},b_{\mathrm{x}}\}$. Similarly, for agent y, the maximum shifting capability is $\min\{r_{\mathrm{y}},b_{\mathrm{y}}\}$ if $b_{\mathrm{y}}>0$, or $\max\{-r_{\mathrm{y}},b_{\mathrm{y}}\}$ if $b_{\mathrm{y}}<0$.
III Equilibrium analysis of the CP shaving game
In this section, we first show the properties of the two-agent two-period CP shaving game $G$ with the specific agent type and prove the pure-strategy NE of the game.
@@ III-A Game properties
We first show the property of CP shaving game $G$. Determined by the parameters of all agents, the game performs differently.
@@
Concave, quasiconcave/discontinuous, and non-concave/discontinuous CP shaving game. Given critical point $r_{\mathrm{x}},r_{\mathrm{y}}$ and balance point $b_{\mathrm{x}},b_{\mathrm{y}},b$ as define in Definition 2, the two-agent two-period CP shaving game $G$ satisfies one of the following:
@@
    Concave CP shaving game. The game $G$ is concave if all agents’ payoff functions $f_{\mathrm{x}}(x,y)$ ($f_{\mathrm{y}}(x,y)$) are concave in $x\in\mathcal{X}_{\mathrm{x}}$ ($y\in\mathcal{X}_{\mathrm{y}}$) for each $y\in\mathcal{X}_{\mathrm{y}}$ ($x\in\mathcal{X}_{\mathrm{x}}$). This concavity condition is satisfied if the following holds
    $b$ $>r_{\mathrm{x}}+r_{\mathrm{y}}.$ (6a)
@@
    Quasiconcave CP shaving game. The game $G$ is quasiconcave/discontinuous if all agents’ payoff functions $f_{\mathrm{x}}(x,y)$ ($f_{\mathrm{y}}(x,y)$) are quasiconcave in $x\in\mathcal{X}_{\mathrm{x}}$ ($y\in\mathcal{X}_{\mathrm{y}}$) for each $y\in\mathcal{X}_{\mathrm{y}}$ ($x\in\mathcal{X}_{\mathrm{x}}$[Reny,, 1999]. This concavity condition is satisfied if the following holds
    $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},\ -r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}};$ (6b)
@@
    Non-concave CP shaving game. The game $G$ is non-concave/discontinuous if it is not concave or quasiconcave. The condition is
    $\{0<b\leq r_{\mathrm{x}}+r_{\mathrm{y}}\}\cap$ [$[\{b_{\mathrm{x}}>r_{\mathrm{x}}\}\cup\{b_{\mathrm{y}}<-r_{\mathrm{y}}\}\cup\{b_{\mathrm{y}}>r_{\mathrm{y}}\}].$ (6c)
@@
The payoff function is quadratic when there is only one possible CP period in the gaming process, resulting in a concave and continuous game. The indicator function discontinues the game when the CP period changes. According to the quasiconcavity definition, we separate three cases: 1) Agent x can switch the CP period regardless of whether it occurs in period 1 or 2; 2) Agent x can switch the CP period only if it occurs in period 1, and its payoff function is l.s.c. as defined in Definition 1; 3) Agent x can switch the CP period only if it occurs in period 2, and the function is u.s.c. as defined in Definition 1. We show only the first case is able for both agents’ payoff functions to be quasiconcave. Then the non-concave condition is derived from the complementary set of concave and quasiconcave conditions. The detailed proof is provided in the appendix. ∎
This Proposition shows the game’s properties depend on the relationship between the agent’s critical point and balance point, affected by baseline demand $X_{1},X_{2},Y_{1},Y_{2}$, CP price $\pi$, and shifting penalty parameters $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$. Under concave game conditions, the CP period always aligns with the baseline since agents’ demand shifting is insufficient to alter it. If all agents are capable, as defined in Definition 3, the CP shaving game is quasiconcave. Otherwise, if both agents’ demand shifting is sufficient to interchange the CP period but at least one agent is non-capable, the game becomes non-concave. In practical systems, due to the variant agents’ shifting capability, all three types of games are likely to appear. Empirical studies on demand response indicate that some industrial customers with automation can shift demand efficiently, while some are inefficient in shifting [Siano,, 2014]. With these properties of CP shaving game $G$, we then analyze its NE.